The length of the base of the triangle is three times the height. The area of the triangle is 24in squared. Find the length of the base of the triangle.
Answer by oscargut(686)

Let name the length of the base L
and height = H
then L=3H
Area=LH/2=3H(H)/2=3H^2/2=24
then H^2=16
then H=4
and L=3H=12
Answer: length of the base is 12 inches